Now suppose the Central Bank Authority on Country X is concerned about the Inflation that government's expansionary fiscal policy may yield, and so it raises the Reserve Requirements from 7% to 15%.

(a) If the total deposits on a month are equal to $200mi, what is the value of Excess Reserves under the new requirements?

(b) How much money could the banking system potentially create?
Answer #1

a) total deposits are equal to $200 million. Reserve requirement is 15%. Therefore the amount of reserves will be 15% of 200 or $30 million. Required reserves are $30 million and the remaining $170 million are excess reserves

b) money multiplier is 1 / required reserves ratio = 1/15% = 6.67. total money created in the banking system = 200*6.67 = $1333.33 million